# Benchmarking ERP Analysis: Manual Features, Deep Learning, and Foundation Models

### v1

@@ Abstract
Event-related potential (ERP), a specialized paradigm of electroencephalographic (EEG), reflects neurological responses to external stimuli or events, generally associated with the brain’s processing of specific cognitive tasks. ERP plays a critical role in cognitive analysis, the detection of neurological diseases, and the assessment of psychological states. Recent years have seen substantial advances in deep learning-based methods for spontaneous EEG and other non-time-locked task-related EEG signals. However, their effectiveness on ERP data remains underexplored, and many existing ERP studies still rely heavily on manually extracted features. In this paper, we conduct a comprehensive benchmark study that systematically compares traditional manual features (followed by a linear classifier), deep learning models, and pre-trained EEG foundation models for ERP analysis. We establish a unified data preprocessing and training pipeline and evaluate these approaches on two representative tasks, ERP stimulus classification and ERP- based brain disease detection, across 12 publicly available datasets. Furthermore, we investigate various patch-embedding strategies within advanced Transformer architectures to identify embedding designs that better suit ERP data. Our study provides a landmark framework to guide method selection and tailored model design for future ERP analysis. The code is available at https://github.com/DL4mHealth/ERP-Benchmark
@@ 1 Introduction
Event-related potential (ERP) is an EEG-derived measure of time-locked neural activity, serving as a foundational tool with broad applications in neuroscience, dementia, and cognitive science [1, 2, 3]. As a specialized category within the broader EEG signal domain, ERP exhibits distinct characteristics compared to spontaneous EEG activity [4]. While spontaneous EEG reflects ongoing intrinsic brain states but with limited functional specificity, ERP captures time-locked neurological responses elicited by specific sensory, audio, or cognitive events [5, 6, 7]. By aligning neural signals to controlled stimuli, ERP provides a direct and interpretable window into transient cognitive processes. The primary advantage of ERP lies in its temporal synchrony between external events and internal brain electrical responses, which typically yields a higher signal-to-noise ratio than spontaneous EEG activity [8]. In practice, ERP trials are extracted from continuous EEG recordings by segmenting event-centered epochs [3]. A short pre-stimulus interval (typically 0.2-0.5 seconds) is used for baseline correction to suppress background noise, followed by a longer post-stimulus interval (typically 0.8-1.5 seconds) that contains the stimulus-evoked neural potential [9]. Figure 1 illustrates a signal comparison between ERP and spontaneous EEG.
@@
In recent years, deep learning has achieved substantial progress in EEG decoding across a wide range of applications, including motor imagery control [10], emotion recognition [11], sleep stage classification [11], seizure detection [12], and dementia detection [13]. Advanced deep learning techniques, such as foundation models that have demonstrated remarkable success in natural language processing, computer vision, and multimodal learning (e.g., ChatGPT [14], Gemini [15], and DINOv2 [16]), are also beginning to be explored for learning generalizable EEG representations. Recent EEG foundation models, including LaBraM [17], EEGPT [18], and CBraMod [19], have reported strong performance on various EEG downstream tasks. However, most existing deep learning model designs primarily target spontaneous EEG signals, such as resting-state and sleep EEG [11], or non-time-locked task-induced EEG, such as motor imagery [20]. In contrast, time-locked evoked EEG signals, exemplified by ERP, remain significantly underexplored. As a result, ERP analysis continues to rely mainly on manual feature extraction and conventional statistical approaches, or directly employs models designed for spontaneous EEG to ERP analysis, resulting in suboptimal performance. Given the distinct signal characteristics of ERP and its fundamental differences from spontaneous EEG, it remains unclear to what extent recent advances in deep learning are directly applicable to ERP decoding. This gap necessitates systematic benchmarking of methods for ERP analysis.
@@
To this end, we conduct a comprehensive and fair benchmark evaluation of ERP classification methods. We implement a wide range of approaches, including 2 manual feature extraction methods (with 31 and 91 handcrafted features, respectively), 10 state-of-the-art deep learning methods for EEG, and 3 advanced EEG foundation models, within a unified data preprocessing and evaluation framework. The benchmark covers two representative and practical ERP classification tasks across 12 datasets. The first focuses on ERP stimulus classification, such as distinguishing standard and target stimuli in oddball paradigms, to evaluate models’ ability to capture fundamental stimulus-locked cognitive responses. The second addresses ERP-based brain disease detection, such as identifying neurological or psychiatric conditions, thereby assessing model utility in more complex and clinically relevant scenarios. Furthermore, to guide the development of more effective models, we conduct a comparative analysis of three prevalent patch-embedding strategies for EEG data within a state-of-the-art Transformer framework. This embedding comparison aims to identify the optimal embedding method for ERP classification, offering critical insights for designing dedicated Transformer backbones for ERP analysis. All the tasks are performed under a rigorously subject-independent setup to evaluate cross-subject learning ability.
@@
Q4: Which patch embedding strategy is most suitable for designing ERP-specific Transformer models?
@@
A4: Uni-variate patch embedding (See Figure 5) achieves better overall performance than multi-variate and whole-variate patching strategies for ERP-specific Transformer.
The rest of this paper is organized as follows: Section 3 presents statistical information about the ERP datasets used in this study, along with the unified preprocessing pipeline. Section 4 describes the methods compared in our benchmark, as well as a test Transformer implemented for embedding comparison. Section 5 provides implementation details and reports experimental results with corresponding analysis. Section 6 discusses the limitations of existing methods and outlines directions for future work. Section 7 concludes the paper.
@@ 2 Related Work
Deep Learning Methods: Deep learning methods with diverse architectures are widely used for EEG decoding. EEGNet [40] is a classic lightweight convolutional model, while EEG-Inception [41] extends temporal convolution with depthwise spatial blocks for ERP-based BCIs. AMWCNN [42] integrates wavelet-based multiscale analysis and attention mechanisms to identify discriminative frequency bands and temporal segments. Transformer-based approaches, such as EEG-Transformer [43] and EEGConformer [44], leverage self-attention to capture long-range temporal dependencies and global contextual information in EEG signals. Neuro-BERT [45] introduces masked autoencoding for robust EEG representation learning, LGGNet [46] explicitly models local–global brain functional graphs, MOCNN [47] proposes multiscale octave convolution for ERP classification, and EEGMamba [48] employs bidirectional Mamba modules for efficient long-sequence EEG modeling.
Foundation Models: Foundation models, which have achieved remarkable success in computer vision and natural language processing domains, are increasingly being explored for EEG representation learning. LaBraM [17] introduces a large-scale EEG foundation model that employs a neural tokenizer to reconstruct spectral representations during pre-training, enabling effective transfer to downstream EEG classification tasks. EEGPT [18] proposes a masked self-supervised learning framework based on high signal-to-noise EEG representations, enhanced by spatio-temporal representation alignment. FORMED [49] repurposes a general time series foundation model to EEG downstream tasks. NeuroLM [50] conceptualizes EEG signals as a foreign language and adopts language-model-inspired architectures to learn unified representations across cognitive and clinical tasks. LUNA [51] presents a topology-agnostic EEG foundation model using cross-attention mechanisms, supporting robust learning under heterogeneous channel montages. CSBrain [52] explores cross-scale spatiotemporal tokenization with structured sparse attention to model dependencies between local and global brain regions.
@@ 3 Datasets
3.1 Datasets
In light of the limited availability of public ERP datasets, we systematically searched all accessible resources known to the authors, including platforms such as OpenNeuro, FigShare, and PhysioNet. We identified 12 datasets, each with a sufficient sample size (40+ subjects) to ensure robust statistical evaluation, including ADHD-WMRI [53], AOPD [54], CESCA-AOOD, CESCA-VOOD, CESCA-FLANKER [55], mTBI-ODD [56], NSERP-MSIT, NSERP-ODD [57], PD-ODD [58], PD-SIM [58], RLPD [59], and SCPD [60]. These datasets span a broad range of cognitive paradigms and clinical conditions, including attention deficit hyperactivity disorder (ADHD), Parkinson’s disease (PD), traumatic brain injury (TBI), and aging-related cognitive decline. To ensure fair and consistent benchmarking, all datasets are preprocessed using a unified ERP preprocessing pipeline (Section 3.2).
@@ 3.2 Data Preprocessing Pipeline
To get ERP trials for training, we apply the following unified preprocessing pipeline. 1) Removal of non-EEG channels: All non-EEG channels are removed, such as EOG or coordinate information. 2) Notch and band-pass filtering: A notch filter at 50 Hz or 60 Hz is applied to suppress line noise, followed by a band-pass filter between 0.5 Hz and 45 Hz. 3) Bad channel interpolation: Bad channels are interpolated when marked. 4) Average re-referencing: Average re-referencing is applied to reduce global noise and potential baseline shifts. 5) Artifact removal: Independent component analysis (ICA), combined with ICLabel [61], is used to automatically identify and remove components associated with eye blinks, muscle activity, and cardiac artifacts. 6) Resampling: All recordings are resampled to a uniform sampling rate of 200 Hz. 7) Baseline correction: Baseline correction is applied to enhance stimulus-related ERP components relative to background noise. 8) Trial epoching: EEG recordings are then segmented into ERP trials based on stimulus events, and the non-ERP recordings are discarded (e.g, resting state). The baseline correction and epoch window length vary across ERP tasks and datasets. 9) Z-Score Normalization: Each channel of every segmented ERP trial is normalized independently to zero mean and unit variance. The statistics of the processed datasets are summarized in Table 1. More details of the preprocessing are provided in our GitHub repository.
@@ 4.1.2 ERP Feature Extraction
Peak-Related Features: amplitudes of positive and negative peaks and their normalized latencies [24, 25, 26]. Peak-related features are extracted to characterize salient ERP components. The maximum positive and minimum negative peak amplitudes reflect the strength of event-related neural responses. The total analysis window length-normalized latencies to account for variability in ERP duration and ensure comparability across trials and subjects.
@@ 4.2.2 ModernTCN
ModernTCN [70] is a convolutional architecture designed for general time-series analysis and demonstrates strong performance, particularly in classification. It combines depthwise convolutions with multiple pointwise convolutions and introduces channel-independent embedding mechanisms. This design effectively addresses key limitations of convolutional models in multivariate time-series tasks, where performance is often constrained by small receptive fields and difficulties in modeling inter-variable dependencies.
@@ 4.2.4 PatchTST
The iTransformer [73] proposes a novel data-embedding method based on the transformer architecture for multivariate time-series analysis. Its distinctive feature is the reversal of the conventional tokenization strategy. Instead of using multi-channel samples at a single time point into a single temporal token, iTransformer treats each channel/variate as an independent variable token. This adjustment enables the self-attention module to model correlations between variables directly, thereby enhancing interpretability.

### v2

@@ Abstract
Event-related potential (ERP), a specialized paradigm of electroencephalographic (EEG), reflects neurological responses to external stimuli or events, generally associated with the brain’s processing of specific cognitive tasks. ERP plays a critical role in cognitive analysis, the detection of neurological diseases, and the assessment of psychological states. Recent years have seen substantial advances in deep learning-based methods for spontaneous EEG and other non-time-locked task-related EEG signals. However, their effectiveness on ERP data remains underexplored, and many existing ERP studies still rely heavily on manually extracted features. In this paper, we conduct a comprehensive benchmark study that systematically compares traditional manual features (followed by a linear classifier), deep learning models, and pre-trained EEG foundation models for ERP analysis. We establish a unified data preprocessing and training pipeline and evaluate these approaches on two representative tasks, ERP stimulus classification and ERP-based brain disease detection, across 12 publicly available datasets. Furthermore, we investigate various token-embedding strategies within advanced Transformer architectures to identify embedding designs that better suit ERP data. Our study provides a landmark framework to guide method selection and tailored model design for future ERP analysis. The code is available at https://github.com/DL4mHealth/ERP-Benchmark
@@ 1 Introduction
ERP is an EEG-derived measure of time-locked neural activity, serving as a foundational tool with broad applications in neuroscience, dementia, and cognitive science [1, 2, 3]. As a specialized category within the broader EEG signal domain, ERP exhibits distinct characteristics compared to spontaneous EEG activity [4]. While spontaneous EEG reflects ongoing intrinsic brain states but with limited functional specificity, ERP captures time-locked neurological responses elicited by specific sensory, audio, or cognitive events [5, 6, 7]. By aligning neural signals to controlled stimuli, ERP provides a direct and interpretable window into transient cognitive processes. The primary advantage of ERP lies in its temporal synchrony between external events and internal brain electrical responses, which typically yields a higher signal-to-noise ratio than spontaneous EEG activity [8]. In practice, ERP trials are extracted from continuous EEG recordings by segmenting event-centered epochs [3]. A short pre-stimulus interval (typically 0.2-0.5 seconds) is used for baseline correction to suppress background noise, followed by a longer post-stimulus interval (typically 0.8-1.5 seconds) that contains the stimulus-evoked neural potential [9]. Figure 1 illustrates a signal comparison between ERP and spontaneous EEG.
@@
In recent years, deep learning has achieved substantial progress in EEG decoding across a wide range of applications, including motor imagery control [10], emotion recognition [11], sleep stage classification [11], seizure detection [12], and dementia detection [13]. In addition, advanced deep learning techniques, such as foundation models that have demonstrated remarkable success in domains like natural language processing (NLP) and computer vision (CV) (e.g., ChatGPT [14], Gemini [15], and DINOv2 [16]), are also being explored for learning generalizable EEG representations. Recent EEG foundation models, including LaBraM [17], EEGPT [18], and CBraMod [19], have reported strong performance on various EEG downstream tasks. However, most existing deep learning model designs primarily target spontaneous EEG signals, such as resting-state and sleep EEG [11], or non-time-locked task-induced EEG, such as motor imagery [20]. In contrast, time-locked evoked EEG signals, exemplified by ERP, remain significantly underexplored. As a result, ERP analysis continues to rely mainly on manual feature extraction and conventional statistical approaches, or directly applies models designed for spontaneous EEG on ERP analysis, resulting in suboptimal performance. Given the distinct signal characteristics of ERP and its fundamental differences from spontaneous EEG, it remains unclear to what extent recent advances in deep learning are directly applicable to ERP decoding. This gap necessitates systematic benchmarking of methods for ERP analysis.
@@
To this end, we conduct a comprehensive and fair benchmark evaluation of ERP classification methods. We implement a wide range of approaches, including 2 manual feature extraction methods (with 31 and 91 handcrafted features, respectively), 10 state-of-the-art deep learning methods for EEG, and 3 advanced EEG foundation models, within a unified data preprocessing and evaluation framework. The benchmark covers two representative and practical ERP classification tasks across 12 public datasets. The first focuses on ERP stimulus classification, such as distinguishing standard and target stimuli in oddball paradigms, to evaluate models’ ability to capture fundamental stimulus-locked cognitive responses. The second addresses ERP-based brain disease detection, such as identifying neurological or psychiatric conditions, thereby assessing model utility in more complex and clinically relevant scenarios. Furthermore, to guide the future development of more effective models, we conduct a comparative analysis of three prevalent token-embedding strategies for EEG data used in recent Transformer architectures. This embedding comparison aims to identify the optimal embedding method for ERP classification, offering critical insights for designing dedicated Transformer backbones for ERP analysis. All the tasks are performed under a rigorously subject-independent setup to evaluate cross-subject learning ability.
@@
Q4: Which token embedding strategy is most suitable for designing ERP-specific Transformer models?
@@
A4: Uni-variate token embedding (See Figure 3) achieves better overall performance than multi-variate and whole-variate patching strategies for ERP-specific Transformer.
@@ 2 Related Work
Deep Learning Methods: Deep learning methods with diverse architectures are widely used for EEG decoding. EEGNet [40] is a classic lightweight convolutional model, while EEG-Inception [41] extends temporal convolution with depthwise spatial blocks for ERP-based BCIs. Transformer-based approaches, such as EEG-Transformer [42] and EEGConformer [43], leverage self-attention to capture long-range temporal dependencies and global contextual information in EEG signals. Neuro-BERT [44] introduces masked autoencoding for robust EEG representation learning, LGGNet [45] explicitly models local-global brain functional graphs, MOCNN [46] proposes multiscale octave convolution for ERP classification, and EEGMamba [47] employs bidirectional Mamba modules for efficient long-sequence EEG modeling.
Foundation Models: Foundation models, which have achieved remarkable success in CV and NLP domains, are increasingly being explored for EEG representation learning. LaBraM [17] introduces a large-scale EEG foundation model that employs a neural tokenizer to reconstruct spectral representations during pre-training. EEGPT [18] proposes a masked self-supervised learning framework based on high signal-to-noise EEG representations, enhanced by spatio-temporal representation alignment. FORMED [48] repurposes a general time series foundation model to EEG downstream tasks. NeuroLM [49] conceptualizes EEG signals as a foreign language and adopts language-model-inspired architectures to learn unified EEG representations. LUNA [50] presents a topology-agnostic EEG foundation model using cross-attention mechanisms, supporting robust learning under heterogeneous channel montages. CSBrain [51] explores cross-scale spatiotemporal tokenization with structured sparse attention to model dependencies between local and global brain regions.
@@ 3 Datasets
3.1 Datasets Selection
In light of the limited availability of public ERP datasets, we systematically searched all accessible resources known to the authors, including platforms such as OpenNeuro, FigShare, and PhysioNet. We identified 12 datasets, each with a sufficient subject size (40+ subjects) to ensure robust statistical evaluation, including ADHD-WMRI [52], AOPD [53], CESCA-AOOD, CESCA-VOOD, CESCA-FLANKER [54], mTBI-ODD [55], NSERP-MSIT, NSERP-ODD [56], PD-ODD [57], PD-SIM [57], RLPD [58], and SCPD [59]. These datasets span a broad range of ERP paradigms, such as oddball (ODD), Simon conflict (SIM), and flanker, as well as various clinical conditions, including attention deficit hyperactivity disorder (ADHD), Parkinson’s disease (PD), traumatic brain injury (TBI), and aging-related cognitive decline. To ensure fair and consistent benchmarking, all datasets are preprocessed using a unified ERP preprocessing pipeline.
@@ 3.2 Data Preprocessing Pipeline
To get ERP trials for training, we apply the following unified preprocessing pipeline (Figure 2). 1) Removal of non-EEG channels: All non-EEG channels are removed, such as EOG or coordinate information. 2) Notch and band-pass filtering: A notch filter at 50 Hz or 60 Hz is applied to suppress line noise, followed by a band-pass filter between 0.5 Hz and 45 Hz. 3) Bad channel interpolation: Bad channels are interpolated when marked. 4) Average re-referencing: Average re-referencing is applied to reduce global noise and potential baseline shifts. 5) Artifact removal: Independent component analysis (ICA), combined with ICLabel [60], is used to automatically identify and remove components associated with eye blinks, muscle activity, and cardiac artifacts. 6) Resampling: All recordings are resampled to a uniform sampling rate of 200 Hz. 7) Baseline correction: Baseline correction is applied to enhance stimulus-related ERP components relative to background noise. 8) Trial epoching: EEG recordings are then segmented into ERP trials based on stimulus events, and the non-ERP recordings are discarded (e.g, resting state). The baseline correction and epoch window length vary across ERP tasks and datasets. 9) Z-Score Normalization: Each channel of every segmented ERP trial is normalized independently to zero mean and unit variance. The statistics of the processed datasets are summarized in Table 1. More details of per-dataset preprocessing are provided in our GitHub repository.
@@ 4 Method
| Methods | Parameters (Million) | Backbone | Pre-training | Method Innovations |
| TCN | 1.1 | CNN | No | Uses 1D causal dilated convolutions to enlarge the receptive field. |
| ModernTCN | 1.2 – 13 | CNN | No | Combines depthwise-pointwise convolutions with channel-independent embeddings to capture long-range and cross-variable dependencies. |
| TimesNet | 2.3 | CNN | No | Transforms 1D time-series into multi-periodicity 2D representations and applies 2D-CNN to capture both intra- and inter-period temporal variations. |
| PatchTST | 0.8 – 2.5 | Transformer | No | Introduces single-channel patch embedding and channel-independent modeling for time-series Transformers. |
| iTransformer | 0.8 – 0.9 | Transformer | No | Treats each variable as an independent token, enabling self-attention to model inter-variable dependencies in multivariate time series. |
| Medformer | 2.4 – 4.7 | Transformer | No | A Transformer model for medical time-series that uses cross-channel patching, multi-granularity embeddings, and hierarchical self-attention. |
| MedGNN | 4.2 – 4.7 | GNN & Transformer | No | A multi-scale graph model that uses adaptive graphs and frequency-domain convolution and graph transformers for medical time-series learning. |
| EEGNet | 0.002 | CNN | No | A compact CNN architecture that uses temporal, depthwise, and separable convolutions to efficiently capture temporal and spatial features in EEG. |
| EEGInception | 0.8 | CNN | No | Uses Inception-style multi-scale temporal convolutions with separable convolutions to capture EEG features across multiple temporal scales. |
| EEGConformer | 1.1 – 2.9 | CNN & Transformer | No | Combines convolutional spatial-temporal embedding with Transformer self-attention to capture local EEG patterns and global dependencies. |
| BIOT | 3.2 | Transformer | Yes | A foundation model for biomedical signals that single-channel patch embeddings with channel tags and self-attention to capture temporal and spatial dependencies. Pre-trained on 2 EEG datasets and 1 ECG dataset. |
| LaBraM | 5.8 | Transformer | Yes | An EEG foundation model that learns EEG representations via vector-quantized neural tokenization and masked EEG reconstruction with a Transformer. Pre-trained on 16 datasets with 2500+ hours of EEG. |
| CBraMod | 8.1 – 14 | Transformer | Yes | An EEG foundation model using a criss-cross Transformer with parallel spatial and temporal attention to model channel and temporal dependencies. Pre-trained on the TUEG dataset with 9000+ hours of EEG. |
@@ 4.1.2 ERP Feature Extraction
Peak-Related Features: amplitudes of positive and negative peaks and their normalized latencies [24, 25, 26]. Peak-related features are extracted to characterize salient ERP components. The maximum positive and minimum negative peak amplitudes reflect the strength of event-related neural responses. The total analysis window-normalized latencies to account for ERP duration variability and ensure comparability across trials.
@@ 4.2.2 ModernTCN
ModernTCN [69] is a convolutional architecture designed for general time-series analysis and demonstrates strong performance, particularly in classification. It combines depthwise convolutions with multiple pointwise convolutions and introduces channel-independent embedding mechanisms. This design effectively addresses key limitations of convolutional models in multi-variate time-series tasks, where performance is often constrained by small receptive fields and difficulties in modeling inter-variable dependencies.
@@ 4.2.4 PatchTST
The iTransformer [72] proposes a novel data-embedding method based on the transformer architecture for multi-variate time-series analysis. Its distinctive feature is the reversal of the conventional tokenization strategy. Instead of using multi-channel samples at a single time point into a single temporal token, iTransformer treats each channel/variate as an independent variable token. This adjustment enables the self-attention module to model correlations between variables directly, thereby enhancing interpretability.
@@ 4.2.10 EEGConformer
4.3 Foundation Models
4.3.1 BIOT
4.3.2 LaBraM
LaBraM [17] is a foundation model for learning universal EEG representations through large-scale unsupervised pre-training. The framework’s core lies in a vector-quantization neural spectral prediction method: First, a neural tokenizer is trained to encode continuous EEG segments into a compact neural lexicon by reconstructing the Fourier spectrum of the raw EEG signals. Subsequently, a neural Transformer is pre-trained using masked EEG modeling (randomly masking portions of EEG segments and predicting their corresponding neural lexical entries) based on this lexicon. Equipped with learnable temporal and spatial embeddings, this model adaptively learns from EEG inputs of arbitrary channel counts and durations. This method uses more than 2,500 hours of EEG data for pre-training. We load their pre-trained weights for fine-tuning.
@@ 4.3.2 LaBraM
4.3.3 CBraMod
4.4 Token Embedding Comparison for ERP Transformer
Existing EEG Transformer models commonly employ three types of token embeddings: multi-variate, uni-variate, and whole-variate. For example, Medformer [73] employs multi-variate patch embedding, PatchTST [71] adopts uni-variate patch embedding, and iTransformer [72] uses whole-variate patch embedding. Given a multi-variate EEG input sample and a patch length , multi-variate embedding uses patches , uni-variate embedding uses patches , and whole-variate embedding uses patches . These patches are then linearly projected into token embeddings that serve as inputs to the Transformer encoder. Figure 3 illustrates the three embedding strategies.
5 Experiments
5.1 Experimental Setups
5.2 ERP Stimulus Classification Results
We first investigate stimulus-type classification in ERP data, with results summarized in Tables 3. All experiments are conducted under a subject-independent evaluation setup, aiming to identify stimulus-related ERP patterns that generalize across subjects. The Top-1, Top-2, and Top-3 performing methods are highlighted in red, blue, and green, respectively. Across the six datasets (CESCA-AOOD, CESCA-VOOD, CESCA-FLANKER, mTBI-ODD, NSERP-MSIT, and NSERP-ODD), the best method achieved F1 scores of 57.18%, 69.64%, 64.33%, 65.79%, 38.51%, and 68.40%, respectively. These results are substantially higher than random performance, indicating that meaningful and consistent stimulus-related ERP patterns can be extracted even under cross-subject evaluation. Among the three CESCA datasets, the visual oddball and flanker paradigms exhibit higher discriminability than the auditory oddball paradigm. Since these datasets share the same group of subjects and recording conditions, this observation suggests that visual oddball and flanker stimuli elicit more salient and distinguishable neural responses than passive auditory oddball stimuli. For the comparison of method performance, we observe that the top-tier results are predominantly achieved by deep learning models trained from scratch. In contrast, existing foundation models for medical time series and EEG do not demonstrate clear performance advantages when fine-tuned from their released checkpoints. Moreover, manually extracted features rarely achieve top-tier performance, and extracted ERP-specific features consistently outperform generic EEG features in this stimulus type classification task.